### v1

@@ 2.3 Stellar velocity dispersion
We derive the stellar velocity dispersion ($\sigma_{*}$) of pressure-dominated galaxies in the simulations. We also note that the stellar velocity dispersion we derived here differs slightly from the velocity dispersion used to compute $v/\sigma$. Following Sohn2024a who derive $\sigma_{*}$ from IllustrisTNG (see also Sohn2022; Sohn2024b), we compute the velocity dispersion of member stellar particles of subhalos within a cylindrical volume that penetrates the center of subhalos. We particularly compute the line-of-sight velocity dispersion along with the $z$-axis, mocking the velocity dispersion measurements based on fiber spectroscopy. We compute the stellar mass-weighted standard deviation as stellar velocity dispersions.
Sohn2024a investigate the velocity dispersions based on various definitions in IllustrisTNG, including different viewing axes and measurement techniques. They show that the 1D velocity dispersions measured along the $x$-, $y$-, and $z$-axes are generally consistent to each other. Furthermore, they derive velocity dispersions based on various techniques, including standard deviation, luminosity- or stellar mass-weighted standard deviation, and bi-weight measurements. All velocity dispersion measurements are consistent with each other. Similarly, the velocity dispersion measurements for EAGLE simulations are also insensitive to the various definitions of velocity dispersion. We thus use the stellar mass-weighted standard deviation of stellar particles along with the $z$-axis as our velocity dispersion measurements (i.e., $\sigma_{*,z}$).
We adopt a fixed 10 kpc aperture to measure stellar velocity dispersions. Using a fixed aperture facilitates comparison with observational measurements, which are typically derived from fiber spectroscopy with fixed angular sizes. Sohn2024b and Sohn2024a used a 3 kpc fixed aperture to study the stellar velocity dispersions of subhalos in the IllustrisTNG-300 simulation. Here, we employ a slightly larger aperture of 10 kpc, which is sufficiently large to mitigate the effects of gravitational softening in the simulations (i.e., $\sim 0.7$ kpc for EAGLE simulations, Schaye2015). Figure 3 shows the relationship between stellar velocity dispersion measured within 10 kpc ($\sigma_{*,10\penalty 10000\ \mathrm{kpc}}$) and stellar mass ($M_{*}$) for subhalos in the EAGLE-50 simulation. For comparison, we also show the same relation using velocity dispersions measured within 3 kpc and $R_{h,*}$ apertures (the solid and the dashed lines in Figure 3). The resulting $M_{*}-\sigma_{*}$ relations are generally consistent with one another, suggesting that the choice of aperture does not significantly affect the subsequent analysis. Hereafter, we refer to $\sigma_{*,z,10\penalty 10000\ \rm kpc}$ as $\sigma_{*}$.
@@ 3.1 The $M_{*}-\sigma_{*}$ Relations
Figure 4: The stellar mass to 1D stellar velocity dispersion measured within a cylindrical volume with a 10 kpc aperture along with $z$-axis ($\sigma_{*,z,10\rm\penalty 10000\ kpc}$) relations of galaxy subhalos in (a) EAGLE-50, (b) EAGLE-eAGN-dT, (c) EAGLE-eAGN-Visc, (d) EAGLE-wAGN, and (e) EAGLE-NoAGN. White circles show the median $\sigma_{*,z,10\rm\penalty 10000\ kpc}$ at each stellar mass bin. Black bold dashed lines show the central 95% of the $\sigma_{*,z,10\rm\penalty 10000\ kpc}$ distributions. Vertical dotted lines indicate the stellar mass limit we applied. Horizontal dotted lines mark the velocity dispersion limit where the stellar mass limit intersects with the 95% upper limit of the velocity dispersion distributions. Below this velocity dispersion limit, our EAGLE galaxy samples are incomplete in terms of velocity dispersion. The subhalos in the shaded region are excluded from the $\sigma_{*}$ complete sample. The blue dashed lines indicate the $M_{*}-\sigma_{*}$ relation for TNG100 (Sohn2024a), and the gray solid lines show the observed $M_{*}-\sigma_{*}$ relation for SDSS quiescent subhalos in the local universe (Zahid2016).
@@
Gray solid lines in Figure 4 display the observed $M_{*}-\sigma_{*}$ relation from Zahid2016. The observed relation is derived from quiescent galaxies with Dn4000 larger than 1.5 selected from SDSS spectroscopic sample within $0.02<z<0.2$. The observed relation is based on $\sigma_{*}$ measured within a 3 kpc aperture. Because the simulated velocity dispersions we compute vary little within 3 kpc and 10 kpc apertures, we compare the observed relation without applying any aperture correction. The observed relation shows a broken power-law with different slopes at $M_{*}=10^{10.26}$ M (see also Nigoche-Netro2011; Cappellari2013a; Cappellari2016).
@@
The $M_{*}-\sigma_{*}$ relation for EAGLE-NoAGN galaxies is much closer to the observed relation compared to the similar relations for galaxies in other simulations (Figure 4). In particular, galaxies with $M_{*}>10^{11}$ M in EAGLE-NoAGN have comparable velocity dispersions with the observed galaxies. This apparent consistency clearly indicates that the AGN feedback has an impact on the velocity dispersion distributions of galaxies. We discuss the impact of the AGN feedback in Section 4.2.
@@
We also compare the EAGLE $M_{*}-\sigma_{*}$ relations with IllustrisTNG-100 (TNG100) simulation. IllustrisTNG simulations use the AGN feedback models implementing the thermal injection (quasar mode) and kinetic wind (kinetic mode) from the AGN (Weinberger2018; Pillepich2018), while EAGLE simulations only use the quasar mode (Schaye2015). We choose TNG100 because the mass resolution of TNG100 is comparable with EAGLE-50 simulations we use. There are TNG50 simulations covering a similar volume with EAGLE-50 simulations, but their mass resolution is an order of magnitude better than the EAGLE-50 simulation. Sohn2024a demonstrated that the $M_{*}-\sigma_{*}$ relations from TNG50 and TNG100 are offset; the stellar velocity dispersion for TNG50 galaxies are slightly larger than that for TNG100 galaxies at given stellar mass.
@@
We use the TNG100 $M_{*}-\sigma_{*,z,3\penalty 10000\ \rm kpc}$ relation derived by Sohn2024a. The TNG100 relation is derived from quiescent galaxies with specific star formation rate lower than $2\times 10^{-11}$ yr-1 at $z=0$. Because Sohn2024a showed that the velocity dispersions measured within 3 kpc and 10 kpc are comparable, we use $M_{*}-\sigma_{*,z,3\penalty 10000\ \rm kpc}$ relation from Sohn2024a without any aperture correction.
@@
In Figure 4, the navy dashed line shows the TNG100 $M_{*}-\sigma_{*,z,3\penalty 10000\ \rm kpc}$ relation. The TNG100 relation describes the distribution of EAGLE-50 quiescent galaxies very well, except for galaxies in the EAGLE-NoAGN simulation. The relations derived from both simulations differ significantly from the observed relation as Sohn2024a already pointed out. In other words, the $M_{*}-\sigma_{*}$ relations derived from EAGLE and TNG100 simulations with AGN feedback are all coincident at $M_{*}>10^{10.5}\penalty 10000\ \rm{M}_{\odot}$.
@@ 3.2 The Velocity Dispersion Complete Samples
The velocity dispersion function (VDF) measures the number of galaxies within bins of velocity dispersion. Deriving the VDF in both observations and simulations can be affected by systematic incompleteness introduced by sample selection. For example, Sohn2017b showed that constructing the VDF from a volume-limited sample leads to incompleteness in velocity dispersion, primarily due to the large scatter between velocity dispersion and galaxy magnitude. Similarly, deriving VDFs from simulations also requires care in constructing velocity dispersion-complete samples (Sohn2024b).
We start constructing the velocity dispersion complete samples from the stellar mass-limited samples (i.e., $M_{*}>10^{9}\penalty 10000\ \rm{M}_{\odot}$) in five EAGLE simulations. Figure 4 displays that stellar velocity dispersions are generally proportional to the stellar mass, but with significant scatters. Because of this scatter, the determination of the velocity dispersion completeness limit of the stellar-mass limited sample is not straightforward.
Following Sohn2024b, we determine the velocity dispersion completeness limit of the EAGLE stellar mass limited samples (see also Sohn2017b). We first derive the velocity dispersion boundaries (black dashed lines in Figure 4) where 95% of galaxies in each stellar mass bin are included. We then compute the intercept between the upper 95% boundary and the stellar mass limit (i.e., $M_{*}=10^{9}\penalty 10000\ \rm{M}_{\odot}$). This intercept (horizontal lines in Figure 4) indicates the velocity dispersion limit of the stellar mass-limited sample. In other words, the galaxies with stellar mass lower than the stellar mass limit may have larger velocity dispersion than the velocity dispersion limit. Thus, constructing the VDF based on the stellar mass-limited sample may cause the incompleteness of the VDF below the velocity dispersion completeness limit.
The velocity dispersion completeness limits are $\log\sigma_{*}\approx 1.7$ across all five simulations. The velocity dispersion completeness limit for EAGLE simulations is consistent with the completeness limit derived from TNG stellar mass limited sample with $\log(M_{*}/\rm{M}_{\odot})>9$ ($\log\sigma_{*}\approx 1.7$; Sohn2024b). We thus compare the EAGLE and TNG VDFs without concern for differences in completeness. The completeness limits for both EAGLE and TNG are slightly lower than the observed limit ($\log\sigma_{*}\approx 1.84$) derived from SDSS magnitude- and volume-limited samples (Sheth2003; Choi2007; Sohn2017b).
The final velocity dispersion complete samples for five EAGLE simulations contain 550 – 750 pressure-dominated galaxies. We use these velocity dispersion complete samples for our velocity dispersion function measurements in Section 3.3. Table 3 summarizes the number of pressure-dominated galaxy samples in EAGLE simulations we used.
Table 3: The number of pressure-dominated galaxies in EAGLE simulations222
@@
3.3 The Velocity Dispersion Functions in EAGLE simulations
Figure 5: (a) Stellar mass functions and (b) stellar velocity dispersion functions of pressure-dominated galaxies in five EAGLE simulations. Green crosses show the relations from EAGLE-NoAGN. The other symbols are the same in Figure 1. Error bars indicate Poisson errors. For bins with fewer than 10 galaxies, we display only the upper values, marked by arrows.
Figure 5 (a) shows the stellar mass functions (SMFs) of pressure-dominated galaxies in the $\sigma_{*}$ complete samples for the five EAGLE simulations. We normalize the stellar mass function by the total three-dimensional volume of the simulations. Blue triangles, red diamonds, magenta squares, orange hexagons, and green crosses represent the SMFs for EAGLE-50, EAGLE-eAGN-dT, EAGLE-eAGN-Visc, EAGLE-wAGN, and EAGLE-NoAGN, respectively. Error bars represent Poisson uncertainties. We note that some high stellar mass or stellar velocity dispersion bins contain fewer than 10 galaxies. For these bins, we show only the upper limit, defined as the upper end of the $1\sigma$ confidence interval, and omit the lower error bar.
The SMFs from the simulations with normal or enhanced AGN feedback models show consistent stellar mass distributions. The simulations with enhanced AGN feedback models contain slightly fewer massive galaxies as expected, but the difference is insignificant. The SMFs derived from simulations with weak and no AGN feedback are flatter at $10.5<\log(M_{*}/\rm{M}_{\odot})<12$ than the SMFs from simulations with AGN feedback. Massive galaxies with $\log(M_{*}/\rm{M}_{\odot})>11$ are more abundant in simulations with no or weak AGN feedback. This comparison indicates that the AGN feedback in EAGLE simulations indeed suppresses the formation of massive galaxies.
@@ 3.4 Comparison between simulated & observed VDFs
We next compare the observed and simulated VDFs of pressure-dominated galaxies. We use the observed VDF from Sohn2017b constructed from the quiescent galaxies in the SDSS spectroscopy. Sohn2017b select quiescent galaxies with $D_{n}4000>1.5$ within the redshift range $0.03\leq z\leq 0.10$ from SDSS main galaxy sample. They obtain the stellar velocity dispersion measured within a fiducial 3 kpc aperture. Sohn2017b demonstrate the importance of building the velocity dispersion complete sample in measuring VDF; they show that the SDSS VDF measured based on the velocity dispersion complete sample is flatter at $\log\sigma_{*}<2.1$ compared to the SDSS VDF measured based on the magnitude or volume-limited samples (Sheth2003; Choi2007; Chae2010). We therefore adopt the field VDF from Sohn2017b, derived from a velocity-dispersion–complete SDSS sample, as it provides the most complete measurement among existing VDF estimates.
@@
Figure 6: (a) Comparison between observed and simulated VDFs. Black circles show the observed VDF derived from SDSS field galaxies at $0.03\leq z\leq 0.10$ (Sohn2017b). Colored symbols display the simulated VDFs from five EAGLE simulations. Symbols are identical from Figure 5. (b) Same as (a), but the simulated VDFs are shifted by the $\sigma_{*}$ offset determined based on the $M_{*}-\sigma_{*}$ relations. Error bars on the solid lines indicate Poisson errors. For bins with fewer than 10 galaxies, we display only the upper values, marked by arrows.
@@
Figure 6 (a) compares directly the observed SDSS VDF (the black line and circles) with the VDFs measured from five EAGLE simulations. All five simulated VDFs show significant offsets compared to the observed VDF. The VDF from EAGLE-NoAGN has a comparable value at $\log\sigma_{*}>2.4$, but it also deviates from the observed VDF at lower velocity dispersions. The offset between the observed and EAGLE VDFs corresponds to $\Delta\log\sigma_{*}\sim 0.15-0.18$ dex, consistent with the offset in the $M_{*}-\sigma_{*}$ relations. The significant offset between observed and EAGLE VDFs is consistent with the result from the comparison between the observed and TNG VDFs (Sohn2024b).
@@
Figure 6 (b) shows the simulated VDFs after shifting in the $\sigma_{*}$ direction, horizontally. To determine the amount of shift, we calculate the mean velocity-dispersion offset between the observed and simulated $M_{*}-\sigma_{*}$ relations (see Figure 4) for more than 30 galaxies within $11<\log(M_{*}/\rm{M}_{\odot})<12$. We apply this mean offset when shifting the VDFs to preserve their overall shapes. Interestingly, the shapes of the simulated VDFs from the normal and enhanced AGN feedback are consistent with the shape of the observed VDF. In particular, the slope of the simulated VDFs at $\log\sigma_{*}>2.4$ is consistent with the observed VDF. On the contrary, the VDFs derived from weak or no AGN feedback show a much flatter slope at a high velocity dispersion end.

### v2

@@ 2.2 Identifying pressure-dominated subhalos in EAGLE
We derived the stellar velocity dispersion ($\sigma_{*}$) of pressure-dominated galaxies in the simulations. We also note that the stellar velocity dispersion we derived here differs slightly from the velocity dispersion used to compute $v/\sigma$. Following Sohn et al. (2024b), which derived $\sigma_{*}$ from IllustrisTNG (see also Sohn et al., 2022, 2024a), we computed the velocity dispersion of member stellar particles of subhalos within a cylindrical volume that penetrates the center of subhalos. In particular, we computed the line-of-sight velocity dispersion along with the $z$-axis, mocking the velocity dispersion measurements based on fiber spectroscopy. We computed the stellar mass-weighted standard deviation as stellar velocity dispersions.
Sohn et al. (2024b) investigated the velocity dispersions based on various definitions in IllustrisTNG, including different viewing axes and measurement techniques. They showed that the 1D velocity dispersions measured along the $x$-, $y$-, and $z$-axes are generally consistent with each other. Furthermore, they derived velocity dispersions based on various techniques, including standard deviation, luminosity- or stellar mass-weighted standard deviation, and bi-weight measurements. All velocity dispersion measurements are consistent with each other. Similarly, the velocity dispersion measurements for EAGLE simulations are also insensitive to the various definitions of velocity dispersion. We thus used the stellar mass-weighted standard deviation of stellar particles, along with the $z$-axis as our velocity dispersion measurements (i.e., $\sigma_{*,z}$).
We adopted a fixed 10 kpc aperture to measure stellar velocity dispersions. Using a fixed aperture facilitates comparison with observational measurements, which are typically derived from fiber spectroscopy with fixed angular sizes. Sohn et al. (2024a) and Sohn et al. (2024b) used a 3 kpc fixed aperture to study the stellar velocity dispersions of subhalos in the IllustrisTNG-300 simulation. Here, we employed a slightly larger aperture of 10 kpc, which is sufficiently large to mitigate the effects of gravitational softening in the simulations (i.e., $\sim 0.7$ kpc for EAGLE simulations; Schaye et al., 2015). Figure 3 shows the relationship between stellar velocity dispersion measured within 10 kpc ($\sigma_{*,10\penalty 10000\ \mathrm{kpc}}$) and stellar mass ($M_{*}$) for subhalos in the EAGLE-50 simulation. For comparison, we also show the same relation using velocity dispersions measured within 3 kpc and $R_{h,*}$ apertures (the solid and dashed lines in Figure 3). The resulting $M_{*}-\sigma_{*}$ relations are generally consistent with one another, suggesting that the choice of aperture does not significantly affect the subsequent analysis. Hereafter, we refer to $\sigma_{*,z,10\penalty 10000\ \rm kpc}$ as $\sigma_{*}$.
@@ 3.1 The $M_{*}-\sigma_{*}$ Relations
Figure 4: Stellar mass to 1D stellar velocity dispersion measured within a cylindrical volume with a 10 kpc aperture along with $z$-axis ($\sigma_{*,z,10\rm\penalty 10000\ kpc}$) relations of galaxy subhalos in (a) EAGLE-50, (b) EAGLE-eAGN-dT, (c) EAGLE-eAGN-Visc, (d) EAGLE-wAGN, and (e) EAGLE-NoAGN. White circles show the median $\sigma_{*,z,10\rm\penalty 10000\ kpc}$ at each stellar mass bin. Bold, dashed black lines show the central 95% of the $\sigma_{*,z,10\rm\penalty 10000\ kpc}$ distributions. Vertical dotted lines indicate the stellar mass limit we applied. Horizontal dotted lines mark the velocity dispersion limit where the stellar mass limit intersects with the 95% upper limit of the velocity dispersion distributions. Below this velocity dispersion limit, our EAGLE galaxy samples are incomplete in terms of velocity dispersion. The subhalos in the shaded region are excluded from the $\sigma_{*}$ complete sample. The dashed blue lines indicate the $M_{*}-\sigma_{*}$ relation for TNG100 (Sohn et al., 2024b), and the solid gray lines show the observed $M_{*}-\sigma_{*}$ relation for SDSS quiescent subhalos in the local Universe (Zahid et al., 2016).
@@
The solid gray lines in Figure 4 display the observed $M_{*}-\sigma_{*}$ relation from Zahid et al. (2016). The observed relation is derived from quiescent galaxies with Dn4000 larger than 1.5 selected from SDSS spectroscopic sample within $0.02<z<0.2$. The observed relation is based on $\sigma_{*}$ measured within a 3 kpc aperture. Because the simulated velocity dispersions we computed vary little within 3 kpc and 10 kpc apertures, we compared the observed relation without applying any aperture correction. The observed relation shows a broken power law with different slopes at $M_{*}=10^{10.26}$ M (see also Nigoche-Netro et al., 2011; Cappellari et al., 2013; Cappellari, 2016).
@@
The $M_{*}-\sigma_{*}$ relation for EAGLE-NoAGN galaxies is much closer to the observed relation compared to the similar relations for galaxies in other simulations (Figure 4). In particular, galaxies with $M_{*}>10^{11}$ M in EAGLE-NoAGN have velocity dispersions comparable with the observed galaxies. This apparent consistency clearly indicates that the AGN feedback has an impact on the velocity dispersion distributions of galaxies. We discuss the impact of the AGN feedback in Section 4.2.
@@
We also compare the EAGLE $M_{*}-\sigma_{*}$ relations with the IllustrisTNG-100 (TNG100) simulation. IllustrisTNG simulations use the AGN feedback models implementing the thermal injection (quasar mode) and kinetic wind (kinetic mode) from the AGN (Weinberger et al., 2018; Pillepich et al., 2018), while EAGLE simulations only use the quasar mode (Schaye et al., 2015). We chose TNG100 because the mass resolution of TNG100 is comparable with the EAGLE-50 simulations we used. There are TNG50 simulations covering a similar volume with EAGLE-50 simulations, but their mass resolution is an order of magnitude better than the EAGLE-50 simulation. Sohn et al. (2024b) demonstrated that the $M_{*}-\sigma_{*}$ relations from TNG50 and TNG100 are offset; the stellar velocity dispersions for TNG50 galaxies are slightly larger than that for TNG100 galaxies at a given stellar mass.
@@
We used the TNG100 $M_{*}-\sigma_{*,z,3\penalty 10000\ \rm kpc}$ relation derived by Sohn et al. (2024b). The TNG100 relation is derived from quiescent galaxies with specific star formation rates lower than $2\times 10^{-11}$ yr-1 at $z=0$. Because Sohn et al. (2024b) showed that the velocity dispersions measured within 3 kpc and 10 kpc are comparable, we used the $M_{*}-\sigma_{*,z,3\penalty 10000\ \rm kpc}$ relation from Sohn et al. (2024b) without any aperture correction.
@@
In Figure 4, the dashed navy line shows the TNG100 $M_{*}-\sigma_{*,z,3\penalty 10000\ \rm kpc}$ relation. The TNG100 relation describes the distribution of EAGLE-50 quiescent galaxies very well, except for galaxies in the EAGLE-NoAGN simulation. The relations derived from both simulations differ significantly from the observed relation, as Sohn et al. (2024b) already pointed out. In other words, the $M_{*}-\sigma_{*}$ relations derived from EAGLE and TNG100 simulations with AGN feedback are all coincident at $M_{*}>10^{10.5}\penalty 10000\ \rm{M}_{\odot}$.
@@ 3.2 The velocity dispersion complete samples
The VDF measures the number of galaxies within bins of velocity dispersion. Deriving the VDF in both observations and simulations can be affected by systematic incompleteness introduced by sample selection. For example, Sohn et al. (2017b) showed that constructing the VDF from a volume-limited sample leads to incompleteness in velocity dispersion, primarily due to the large scatter between velocity dispersion and galaxy magnitude. Similarly, deriving VDFs from simulations also requires care in the construction of velocity dispersion -complete samples (Sohn et al., 2024a).
We started constructing the velocity dispersion -complete samples from the stellar mass-limited samples (i.e., $M_{*}>10^{9}\penalty 10000\ \rm{M}_{\odot}$) in five EAGLE simulations. Figure 4 shows that stellar velocity dispersions are generally proportional to the stellar mass, but with significant scatters. Because of this scatter, the determination of the velocity dispersion completeness limit of the stellar mass-limited sample is not straightforward.
Following Sohn et al. (2024a), we determined the velocity dispersion completeness limit of the EAGLE stellar mass-limited samples (see also Sohn et al., 2017b). We first derived the velocity dispersion boundaries (dashed black lines in Figure 4), where 95% of galaxies in each stellar mass bin are included. We then computed the intercept between the upper 95% boundary and the stellar mass limit (i.e., $M_{*}=10^{9}\penalty 10000\ \rm{M}_{\odot}$). This intercept (horizontal lines in Figure 4) indicates the velocity dispersion limit of the stellar mass-limited sample. In other words, the galaxies with a stellar mass lower than the stellar mass limit may have larger velocity dispersion than the velocity dispersion limit. Thus, constructing the VDF based on the stellar mass-limited sample may cause the incompleteness of the VDF below the velocity dispersion completeness limit.
The velocity dispersion completeness limits are $\log\sigma_{*}\approx 1.7$ across all five simulations. The velocity dispersion completeness limit for EAGLE simulations is consistent with the completeness limit derived from TNG stellar mass-limited sample with $\log(M_{*}/\rm{M}_{\odot})>9$ ($\log\sigma_{*}\approx 1.7$; Sohn et al., 2024a). We thus compared the EAGLE and TNG VDFs without concern for differences in completeness. The completeness limits for both EAGLE and TNG are slightly lower than the observed limit ($\log\sigma_{*}\approx 1.84$) derived from SDSS magnitude- and volume-limited samples (Sheth et al., 2003; Choi et al., 2007; Sohn et al., 2017b).
The final velocity dispersion complete samples for five EAGLE simulations contain 550 – 750 pressure-dominated galaxies. We used these velocity dispersion -complete samples for our VDF measurements in Section 3.3. Table 3 summarizes the number of pressure-dominated galaxy samples in EAGLE simulations that we used.
Table 3: Number of pressure-dominated galaxies in EAGLE simulations.222
@@
3.3 The VDFs in EAGLE simulations
Figure 5: (a) Stellar mass functions and (b) VDFs of pressure-dominated galaxies in five EAGLE simulations. Green crosses show the relations from EAGLE-NoAGN. The other symbols are the same as in Figure 1. Error bars indicate Poisson errors. For bins with fewer than ten galaxies, we only display the upper values, which are marked by arrows.
Figure 5 (a) shows the stellar mass functions (SMFs) of pressure-dominated galaxies in the $\sigma_{*}$ -complete samples for the five EAGLE simulations. We normalized the SMF by the total three-dimensional volume of the simulations. Blue triangles, red diamonds, magenta squares, orange hexagons, and green crosses represent the SMFs for EAGLE-50, EAGLE-eAGN-dT, EAGLE-eAGN-Visc, EAGLE-wAGN, and EAGLE-NoAGN, respectively. Error bars represent Poisson uncertainties. We note that some high stellar mass or stellar velocity dispersion bins contain fewer than ten galaxies. For these bins, we only show the upper limit, defined as the upper end of the $1\sigma$ confidence interval, and omit the lower error bar.
The SMFs from the simulations with normal or enhanced AGN feedback models show consistent stellar mass distributions. The simulations with enhanced AGN feedback models contain slightly fewer massive galaxies, as expected, but the difference is insignificant. The SMFs derived from simulations with weak and no AGN feedback are flatter at $10.5<\log(M_{*}/\rm{M}_{\odot})<12$ than the SMFs from simulations with AGN feedback. Massive galaxies with $\log(M_{*}/\rm{M}_{\odot})>11$ are more abundant in simulations with no or weak AGN feedback. This comparison indicates that the AGN feedback in EAGLE simulations indeed suppresses the formation of massive galaxies.
@@ 3.4 Comparison between simulated & observed VDFs
We next compare the observed and simulated VDFs of pressure-dominated galaxies. We used the observed VDF from Sohn et al. (2017b) constructed from the quiescent galaxies in the SDSS spectroscopy. Sohn et al. (2017b) selected quiescent galaxies with $D_{n}4000>1.5$ within the $0.03\leq z\leq 0.10$ redshift range from the SDSS main galaxy sample. They obtained the stellar velocity dispersion measured within a fiducial 3 kpc aperture. Sohn et al. (2017b) demonstrated the importance of building the velocity dispersion -complete sample for measuring the VDF; they showed that the SDSS VDF measured based on the velocity dispersion -complete sample is flatter at $\log\sigma_{*}<2.1$ compared to the SDSS VDF measured based on the magnitude- or volume-limited samples (Sheth et al., 2003; Choi et al., 2007; Chae, 2010). We therefore adopted the field VDF from Sohn et al. (2017b), derived from a velocity dispersion complete SDSS sample, as it provides the most complete measurement among existing VDF estimates.
@@
Figure 6: (a) Comparison between observed and simulated VDFs. Black circles show the observed VDF derived from SDSS field galaxies at $0.03\leq z\leq 0.10$ (Sohn et al., 2017b). Colored symbols display the simulated VDFs from five EAGLE simulations. Symbols are identical from Figure 5. (b) Same as (a), but the simulated VDFs are shifted by the $\sigma_{*}$ offset determined based on the $M_{*}-\sigma_{*}$ relations. Error bars on the solid lines indicate Poisson errors. For bins with fewer than ten galaxies, we only display the upper values, marked by arrows.
@@
Figure 6 (a) directly compares the observed SDSS VDF (the black line and circles) with the VDFs measured from five EAGLE simulations. All five simulated VDFs show significant offsets compared to the observed VDF. The VDF from EAGLE-NoAGN has a comparable value at $\log\sigma_{*}>2.4$, but it also deviates from the observed VDF at lower velocity dispersions. The offset between the observed and EAGLE VDFs corresponds to $\Delta\log\sigma_{*}\sim 0.15-0.18$ dex, consistent with the offset in the $M_{*}-\sigma_{*}$ relations. The significant offset between observed and EAGLE VDFs is consistent with the result from the comparison between the observed and TNG VDFs (Sohn et al., 2024a).
@@
Figure 6 (b) shows the simulated VDFs after shifting in the $\sigma_{*}$ direction, horizontally. To determine the amount of shift, we calculate the mean velocity dispersion offset between the observed and simulated $M_{*}-\sigma_{*}$ relations (see Figure 4) for more than 30 galaxies within $11<\log(M_{*}/\rm{M}_{\odot})<12$. We applied this mean offset when shifting the VDFs to preserve their overall shapes. Interestingly, the shapes of the simulated VDFs from the normal and enhanced AGN feedback are consistent with the shape of the observed VDF. In particular, the slope of the simulated VDFs at $\log\sigma_{*}>2.4$ is consistent with the observed VDF. On the contrary, the VDFs derived from weak or no AGN feedback show a much flatter slope at a high-velocity dispersion end.